OpenStudy (anonymous):

Expand log: log√(x^7y)/z^5

OpenStudy (anonymous):

\[\log \sqrt{x^{7}y/z^{5}}\]

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

cmon help me out

OpenStudy (anonymous):

\[\log \sqrt{\frac{ x^7y }{ z^5 }}\] is this it?

OpenStudy (anonymous):

yea lol couldnt figure out how to write it that way

OpenStudy (anonymous):

\[\log x^a = a \log x\text{, }\log (u\cdot v)= \log u + \log v\text{, }\log \left( \frac{ u }{ v } \right)= \log u - \log v\]

OpenStudy (anonymous):

use the properties of logs... first rewrite the square root using an exponent

OpenStudy (anonymous):

btw \frac {numerator}{denominator} to write a fraction in the equation editor

OpenStudy (anonymous):

so it would be \[\frac {7\log(x) +\log(y) -5\log(z)}{2}\] ?

OpenStudy (anonymous):

there you go... i'd write it like this: \[\frac{ 7 }{ 2 }\log x +\frac{ 1 }{ 2 }\log y - \frac{ 5 }{ 2 }\log z\] but that's just me. your way is perfectly a okay!

OpenStudy (anonymous):

awesome thanks man

OpenStudy (anonymous):

remember, logs are just exponents so if you remember the properties of exponents, then you already know the properties of logs

OpenStudy (anonymous):

thats true